OpenStudy (anonymous):

HELP I NEED QUICK ANSWER find the circumference of the circle to the nearest tenth. use 3.14

OpenStudy (amistre64):

you will need to use the formula for the circumference of a circle then

OpenStudy (amistre64):

no, im confident enough that i can surmise the specifics of the circle all on my own :/

OpenStudy (amistre64):

is the measure they give a diameter or a radius?

OpenStudy (anonymous):

Do you go to connections academy?

OpenStudy (amistre64):

the goal of this site is to help you study material, as opposed to just giving you an answer to a specific problem. the process involved for this solution is to recall the ratio of pi with the parts of a circle is the measure they give a diameter or a radius?

OpenStudy (amistre64):

the definition of pi is such that\[\pi=\frac{Circumference}{diameter}\] solving for the top part:\[C = d*\pi\]

OpenStudy (amistre64):

in this case, i believe the measure they give is the diameter

OpenStudy (amistre64):

good luck :)